# Learning to Manipulate under Limited Information

### v3

@@ Introduction
As for the difficulty of manipuation, there is now a large literature in computational social choice on the worst-case complexity of computing whether there is a strategic ranking that will elect a desired candidate. Faliszewski and Procaccia (2010) call this “AI’s war on manipulation.” A series of hardness results have been proved since Bartholdi, Tovey, and Trick 1989 and Conitzer, Sandholm, and Lang 2007. There is evidence for and against the view that high worst-case computational complexity is a barrier to manipulation (Walsh 2011), and the situation with average-case manipulability might be quite different (Mossel and Rácz 2013).
@@
A number of previous papers study whether a voter can successfully manipulate in an election under limited information about how other voters will vote (Myerson and Weber 1993; Conitzer, Walsh, and Xia 2011; Reijngoud and Endriss 2012; Meir, Lev, and Rosenschein 2014; Endriss et al. 2016; Lang 2020; Veselova 2023), including with the use of heuristics (Chopra, Pacuit, and Parikh 2004; Laslier 2009; Meir 2018; Fairstein et al. 2019). The information of the manipulator is typically represented by a set of preference profiles, all of which agree on (i) the manipulator’s own preferences and (ii) some other partial information (e.g., all profiles in the manipulator’s information set agree on who will win the election if the manipulator votes sincerely, or all profiles in the set are such that for each of the other voters $i$, $i$’s ranking in the profile extends some known partial order over the candidates that the manipulator attributes to $i$), perhaps supplemented with a probability measure over such profiles (Lu et al. 2012). By contrast, in this paper we will represent limited information by the inputs to a neural network.
@@
We also used a spatial 2D model: each candidate and each voter is independently placed in $\mathbb{R}^{2}$ according to the multivariate normal distribution (as in Merrill 1988) with no correlation between the two dimensions; the utility of a candidate for a voter is the square of the Euclidean distance between the candidate and the voter (using the quadratic proximity utility function as in Merrill and Grofman 1999, p. 21).
@@
These utility profiles can then be parameterized as 2D matrices, $U\in\mathbb{R}^{n\times m}$, where $n$ is the number of voters, $m$ is the number of candidates, and $U[i,c]\in\left[0,1\right]$. To select the utility function for a given voter $i$ is to select the row $U_{i}$.
@@
  • optimizing labeling: all optimal rankings to submit are labeled by $1$, and all other rankings are labeled by 00.777We duplicated our experiments with a satisficing labeling (if there are profitable manipulations, they are labeled by $1$, and all others rankings are labeled by 00; otherwise all rankings that do at least as well as the sincere ranking are labeled by $1$ and all others by 00), but MLPs trained with this labeling had qualitatively indistinguishable results from those trained with the optimizing labeling.
@@
To evaluate how well a given MLP has learned to manipulate, we must convert its output distribution over rankings into a single ranking. To do so, we used the following decision rule:
@@
For each voting method $F$, each $n\in\{5,6,10,11,20,21\}$, each $m\in\{3,4,5,6\}$, each choice of an input type for the MLP, and each choice of a model size (see the $x$-axis of Figure 2), we trained one or more “generations” of MLPs with that model size to manipulate elections with $n$ voters and $m$ candidates run using $F$, resulting in over 100,000 trained MLPs. For a given generation, we used the same initialization of MLP weights and the same training, validation, and evaluation profiles for every MLP for $n$ candidates and $m$ voters. Across generations, we varied the initialization of MLP weights and used different training, validation, and evaluation profiles, to provide reassurance that our results were not due to lucky initial weights or profiles. All elections and labels were pre-computed so training could rely fully on the GPU.
@@
All code was written in Python using PyTorch, version 2.0.1, and the pref_voting library (pypi.org/project/pref-voting/), version 0.4.42 or later. Training and evaluation was parallelized across nine local Apple computers with Apple silicon, the most powerful equipped with an M2 Ultra with 24-core CPU, 76-core GPU, and 128GB of unified memory, running Mac OS 13, as well as up to sixteen cloud instances with Nvidia A6000 or A10 GPUs running Linux Ubuntu 18.04.
@@ Results
Figure 2 and Supplementary Figures A.2-B.3 show the performance of MLPs with each of the 26 different hidden layer configurations we tested, focusing on 6 candidates, 10/11 voters, and different choices of limited information. All Supplementary Figures cited are available at github.com/epacuit/ltm/blob/main/supplementary-figures.pdf.
@@
Figure 2: Results using the uniform utility model with 6 candidates and 10/11 voters for MLPs manipulating on the basis of the plurality scores or majority matrix. Error bars indicate twice the estimated standard error of the mean. Hidden layer configurations of trained MLPs are shown on the x-axis. Versions of this figure for the Mallows model, spatial 2D model, and different types of limited information appear in Supplementary Figures A.2-B.3.
@@
Comparing the results for the three probability models in Figure 1, we see that MLPs achieved the highest profitability in manipulating against voting methods in elections generated by the uniform utility model, followed by the Mallows model, followed by the spatial 2D model. While the graphs for Mallows are roughly scaled down versions of the those for the uniform model, it is striking that for the spatial 2D model, even the best MLP’s could not learn to profitably manipulate against Minimax, Nanson, and Split Cycle. [^12] A natural thought to explain this is that Minimax, Nanson, and Split Cycle are Condorcet consistent, and there is a high frequency of Condorcet winners under the spatial 2D model (yet this must be squared with the results for Stable Voting). However, even when there is a Condorcet winner in $\mathbf{P}$ and we are using a Condorcet voting method, a voter may still have an incentive to submit an insincere ranking in order to create a majority cycle, possibly resulting in a different winner. On the other hand, such possibilities are evidently rare and difficult to learn to exploit. On the other hand, the comparative usefulness, for manipulating against each voting method, of the different types of limited information is largely the same under all three models (this is true even for Minimax, Nanson, and Split Cycle under the spatial model, looking at which types of information produce less negative results). We conjecture that these findings about types of limited information are robust across other standard probability models as well.
@@
With the uniform utility and Mallows models, sufficiently large MLPs learned to profitably manipulate all eight voting methods on the basis of knowing only the majority matrix, though the profitability of such manipulation varied dramatically across methods. Interestingly, we did not find a substantial increase in profitability of manipulation for MLPs that learned to manipulate based on the more informative margin matrix instead of the majority matrix, except in the case of Borda and Black’s (especially for 6 candidates, as shown in Supplementary Figures B.1-B.3). In fact, the qualitative margin matrix was about as useful as the margin matrix for learning to manipulate Borda and Black’s.
@@
Though obviously useful for manipulating Plurality and somewhat useful for several methods for 3 candidates (though less useful for 4 or 5), knowing the plurality scores was insufficient in 6-candidate elections for profitably manipulating methods other than Plurality (though Borda may be barely manipulable in this case), as shown in Figure 2 and Supplementary Figures A.2-A.3. Moreover, in the case of manipulating Plurality, learning to manipulate on the basis of the plurality ranking led to profitability comparable to learning on the basis of the plurality scores themselves (see Figure 1).
@@
The parity of the number of voters is a key factor for some methods. This is most striking for Stable Voting (also see Black’s), which is barely manipulable with 11 voters but more manipulable with 10 (see Figure 2 and Supplementary Figures A.2-A.3). A key difference is that with 10 voters, it is possible to have margins of zero between candidates, in which case a single voter has more manipulative power under Stable Voting, which produces fewer tied elections in the presence of zero margins than other methods like Minimax.
Figure 3: Top: average profitability of submitted rankings by an ideal manipulator. Middle: average profitability by the best performing MLP with any hidden layer configuration using the majority matrix information, averaging over 5, 6, 10, 11, 20, and 21 voters. Bottom: the ratio of the average profitability of the MLP’s submitted ranking to the average profitability of the ideal manipulator’s submitted ranking. Versions of this figure for the Mallows model and the spatial 2D model appear in Supplementary Figures D.2-3.
@@
For an ideal manipulator, manipulation becomes more profitable as the number of candidates increases (see Figure 3 and Supplementary Figures D.2-D.3). The same can be true for the MLP-based manipulators up to a point, e.g., 4 or 5 candidates for most voting methods when using the majority, qualitative margin, or margin information. However, the ratio between the profitability of the MLP-based manipulator’s submitted rankings and those of the ideal manipulator declines as the number of candidates increases from 3 to 6 (again see the figures just cited). This is intuitive, as manipulation should increase in complexity with more candidates.
@@
Not only is it more profitable to manipulate, say, Borda than Stable Voting, but it is also easier to learn—in the sense of requiring smaller MLPs: whether we consider one, two, or three hidden layers, a smaller network is needed to learn to profitably manipulate Borda compared to Stable Voting on the basis of the majority or margin matrix (see Figure 2 and Supplementary Figures A.2-B.3).
@@ Conclusion
In committee-sized elections (5-21 voters), MLPs can learn to vote strategically on the basis of limited information, though the profitability of doing so varies significantly between different voting methods. This serves as a proof of concept for the study of machine learnability of manipulation under limited information. There are a number of natural extensions for future work, including manipulation by a coalition of multiple voters, additional probability models for generating elections, and training an MLP to manipulate in elections with varying numbers of voters and candidates (rather than training different MLPs for each choice of a number of voters and number of candidates). Our code is already set up to handle these extensions, which only require more compute. However, further research is needed on other questions: What if all agents in the election strategize? What is the social cost or benefit of the learned manipulations? How do different neural network architectures affect results? Finally, one limitation of the classification approach in this paper is that it is infeasible to apply to more than 6 candidates. To overcome this limitation, we plan to develop a reinforcement learning approach to learning to manipulate.
@@ Acknowledgments
For helpful feedback, we thank Dominic Hughes, the anonymous reviewers for SCaLA 2024 and for AAAI 2025, and the participants in the Values-Centered AI seminar at the University of Maryland in Fall 2023, the logic seminar at the University of Maryland in Spring 2024, the Interdisciplinary Workshop on Computational Social Choice at the Center for Human-Compatible Artificial Intelligence in Spring 2024, and New Directions in Social Choice at EC 2024.

### v4

@@ Introduction
As for the difficulty of manipulation, there is now a large literature in computational social choice on the worst-case complexity of computing whether there is a strategic ranking that will elect a desired candidate. Faliszewski and Procaccia (2010) call this “AI’s war on manipulation.” A series of hardness results have been proved since Bartholdi, Tovey, and Trick 1989 and Conitzer, Sandholm, and Lang 2007. There is evidence for and against the view that high worst-case computational complexity is a barrier to manipulation (Walsh 2011), and the situation with average-case manipulability might be quite different (Mossel and Rácz 2013).
@@
A number of previous papers study whether a voter can successfully manipulate in an election under limited information about how other voters will vote (Myerson and Weber 1993; Conitzer, Walsh, and Xia 2011; Reijngoud and Endriss 2012; Meir, Lev, and Rosenschein 2014; Endriss et al. 2016; Lang 2020; Veselova 2023), including through the use of heuristics (Chopra, Pacuit, and Parikh 2004; Laslier 2009; Meir 2018; Fairstein et al. 2019). The information of the manipulator is typically represented by a set of preference profiles, all of which agree on (i) the manipulator’s own preferences and (ii) some other partial information (e.g., all profiles in the manipulator’s information set agree on who will win the election if the manipulator votes sincerely, or all profiles in the set are such that for each of the other voters $i$, $i$’s ranking in the profile extends some known partial order over the candidates that the manipulator attributes to $i$), perhaps supplemented with a probability measure over such profiles (Lu et al. 2012). By contrast, in this paper we will represent limited information by the inputs to a neural network.
@@
We also used a spatial 2D model: each candidate and each voter is independently placed in $\mathbb{R}^{2}$ according to the multivariate normal distribution (as in Merrill 1988) with no correlation between the two dimensions; the utility of a candidate for a voter is the negative of the square of the Euclidean distance between the candidate and the voter (using the quadratic proximity utility function as in Merrill and Grofman 1999, p. 21).
@@
These utility profiles can then be parameterized as 2D matrices, $U\in\mathbb{R}^{n\times m}$, where $n$ is the number of voters, $m$ is the number of candidates, and $U[i,c]\in\mathbb{R}$. To select the utility function for a given voter $i$ is to select the row $U_{i}$.
@@
  • optimizing labeling: all optimal rankings to submit are labeled by $1$, and all other rankings are labeled by 00.777We duplicated our experiments with a satisficing labeling (if there are profitable manipulations, they are labeled by $1$, and all other rankings are labeled by 00; otherwise all rankings that do at least as well as the sincere ranking are labeled by $1$ and all others by 00), but MLPs trained with this labeling had qualitatively indistinguishable results from those trained with the optimizing labeling.
@@
To evaluate how well a given MLP has learned to manipulate, we must convert its output distribution over rankings into a single ranking. To do so, we use the following decision rule:
@@
For each voting method $F$, each $n\in\{5,6,10,11,20,21\}$, each $m\in\{3,4,5,6\}$, each choice of an input type for the MLP, and each choice of a model size (see the $x$-axis of Figure 2), we trained one or more “generations” of MLPs with that model size to manipulate elections with $n$ voters and $m$ candidates run using $F$, resulting in over 100,000 trained MLPs. For a given generation, we used the same initialization of MLP weights and the same training, validation, and evaluation profiles for every MLP for $n$ voters and $m$ candidates. Across generations, we varied the initialization of MLP weights and used different training, validation, and evaluation profiles, to provide reassurance that our results were not due to lucky initial weights or profiles. All elections and labels were pre-computed so training could rely fully on the GPU.
@@
All code was written in Python using PyTorch, version 2.0.1, and the pref_voting library (pypi.org/project/pref-voting/), version 0.4.42 or later. Training and evaluation were parallelized across nine local Apple computers with Apple silicon, the most powerful equipped with an M2 Ultra with 24-core CPU, 76-core GPU, and 128GB of unified memory, running macOS 13, as well as up to sixteen cloud instances with Nvidia A6000 or A10 GPUs running Linux Ubuntu 18.04.
@@ Results
Figure 2 and Supplementary Figures A.2–B.3 show the performance of MLPs with each of the 26 different hidden layer configurations we tested, focusing on 6 candidates, 10/11 voters, and different choices of limited information. All Supplementary Figures cited are available at github.com/epacuit/ltm/blob/main/supplementary-figures.pdf.
@@
Figure 2: Results using the uniform utility model with 6 candidates and 10/11 voters for MLPs manipulating on the basis of the plurality scores or majority matrix. Error bars indicate twice the estimated standard error of the mean. Hidden layer configurations of trained MLPs are shown on the x-axis. Versions of this figure for the Mallows model, spatial 2D model, and different types of limited information appear in Supplementary Figures A.2–B.3.
@@
Comparing the results for the three probability models in Figure 1, we see that MLPs achieved the highest profitability in manipulating against voting methods in elections generated by the uniform utility model, followed by the Mallows model, followed by the spatial 2D model. While the graphs for Mallows are roughly scaled down versions of those for the uniform model, it is striking that for the spatial 2D model, even the best MLPs could not learn to profitably manipulate against Minimax, Nanson, and Split Cycle. [^12] A natural thought to explain this is that Minimax, Nanson, and Split Cycle are Condorcet consistent, and there is a high frequency of Condorcet winners under the spatial 2D model (yet this must be squared with the results for Stable Voting). However, even when there is a Condorcet winner in $\mathbf{P}$ and we are using a Condorcet voting method, a voter may still have an incentive to submit an insincere ranking in order to create a majority cycle, possibly resulting in a different winner. On the other hand, such possibilities are evidently rare and difficult to learn to exploit. On the other hand, the comparative usefulness, for manipulating against each voting method, of the different types of limited information is largely the same under all three models (this is true even for Minimax, Nanson, and Split Cycle under the spatial model, looking at which types of information produce less negative results). We conjecture that these findings about types of limited information are robust across other standard probability models as well.
@@
With the uniform utility and Mallows models, sufficiently large MLPs learned to profitably manipulate all eight voting methods on the basis of knowing only the majority matrix, though the profitability of such manipulation varied dramatically across methods. Interestingly, we did not find a substantial increase in profitability of manipulation for MLPs that learned to manipulate based on the more informative margin matrix instead of the majority matrix, except in the case of Borda and Black’s (especially for 6 candidates, as shown in Supplementary Figures B.1–B.3). In fact, the qualitative margin matrix was about as useful as the margin matrix for learning to manipulate Borda and Black’s.
@@
While knowing the plurality scores is obviously useful for manipulating Plurality and somewhat useful for several methods for 3 candidates (though less useful for 4 or 5), it was insufficient in 6-candidate elections for profitably manipulating methods other than Plurality (though Borda may be barely manipulable in this case), as shown in Figure 2 and Supplementary Figures A.2–A.3. Moreover, in the case of manipulating Plurality, learning to manipulate on the basis of the plurality ranking led to profitability comparable to learning on the basis of the plurality scores themselves (see Figure 1).
@@
The parity of the number of voters is a key factor for some methods. This is most striking for Stable Voting (also see Black’s), which is barely manipulable with 11 voters but more manipulable with 10 (see Figure 2 and Supplementary Figures A.2–A.3). A key difference is that with 10 voters, it is possible to have margins of zero between candidates, in which case a single voter has more manipulative power under Stable Voting, which produces fewer tied elections in the presence of zero margins than other methods like Minimax.
Figure 3: Top: average profitability of submitted rankings by an ideal manipulator. Middle: average profitability by the best performing MLP with any hidden layer configuration using the majority matrix information, averaging over 5, 6, 10, 11, 20, and 21 voters. Bottom: the ratio of the average profitability of the MLP’s submitted ranking to the average profitability of the ideal manipulator’s submitted ranking. Versions of this figure for the Mallows model and the spatial 2D model appear in Supplementary Figures D.2–3.
@@
For an ideal manipulator, manipulation becomes more profitable as the number of candidates increases (see Figure 3 and Supplementary Figures D.2–D.3). The same can be true for an MLP-based manipulator up to a point, e.g., 4 or 5 candidates for most voting methods when using the majority, qualitative margin, or margin information. However, the ratio between the profitability of the MLP-based manipulator’s submitted rankings and those of the ideal manipulator declines as the number of candidates increases from 3 to 6 (again see the figures just cited). This is intuitive, as manipulation should increase in complexity with more candidates.
@@
Not only is it more profitable to manipulate, say, Borda than Stable Voting, but it is also easier to learn—in the sense of requiring smaller MLPs: whether we consider one, two, or three hidden layers, a smaller network is needed to learn to profitably manipulate Borda compared to Stable Voting on the basis of the majority or margin matrix (see Figure 2 and Supplementary Figures A.2–B.3).
@@ Conclusion
In committee-sized elections (5-21 voters), MLPs can learn to vote strategically on the basis of limited information, though the profitability of doing so varies significantly between different voting methods. This serves as a proof of concept for the study of machine learnability of manipulation under limited information. There are a number of natural extensions for future work, including manipulation by a coalition of voters, additional probability models for generating elections, and training an MLP to manipulate in elections with varying numbers of voters and candidates (rather than training different MLPs for each choice of a number of voters and number of candidates). Our code is already set up to handle these extensions, which only require more compute. However, further research is needed on other questions: What if all agents in the election strategize? What is the social cost or benefit of the learned manipulations? How do different neural network architectures affect results? Finally, one limitation of the classification approach in this paper is that it is infeasible to apply to more than 6 candidates. To overcome this limitation, we plan to develop a reinforcement learning approach to learning to manipulate.
@@ Acknowledgments
For helpful feedback, we thank Dominic Hughes, the anonymous reviewers for SCaLA and for AAAI, and the participants in the Values-Centered AI seminar at the University of Maryland in Fall 2023, the logic seminar at the University of Maryland in Spring 2024, the Interdisciplinary Workshop on Computational Social Choice at the Center for Human-Compatible Artificial Intelligence in Spring 2024, and New Directions in Social Choice at EC 2024.